The probability of an event that must occur is 2. State whether the statement is true...

  1. The probability of an event that must occur is

2. State whether the statement is true or false

Every probability is a number between 0 and 1 inclusive

  

3.The result of an experiment is called an

4. In a sample of 50,000 first-born babies, 5 were found to have Prader-Willi syndrome. Find the probability that a family's first child will be born with this syndrome.

  1. The probability of an event that cannot occur is

QUESTION 6

  1. The sum of the probabilities of all possible outcomes of an experiment is

QUESTION 7

  1. For any event A , P(A) +P(not A) =1

    If the probability that an event occurs is 0.8 , the probability the event does not occur is

QUESTION 8

  1. Any particular outcome or a group of outcomes is an

QUESTION 9

  1. The set of all possible simple events is the
Answer #1

Solution :

1. The probability of an event that must occur is 1.

2.

True, Every probability is a number between 0 and 1 inclusive.

3.

Outcome

4.

probability that a family's first child will be born with Prader-Willi syndrome = 5/50000= 0.0001

5.

The probability of an event that cannot occur is 0.

6.

The sum of the probabilities of all possible outcomes of an experiment is 1.

7.

probability that event doesn't occur is =1-0.8= 0.2

8.

Any particular outcome or a group of outcomes is an sample space.

9.

The set of all possible simple events is the experiment.

​​​​